Maricar is ten times older than her son Carl. In five years, the product of her age and her son's age two years ago will be 35. Find their present age.
Answer by ankor@dixie-net.com(22740)

Maricar is ten times older than her son Carl.
m = 10c
:
In five years, the product of her age and her son's age two years ago will be 35.
(m+5)(c-2) = 35
replace m with 10c
(10c+5)(c-2) = 35
FOIL
10c^2 - 20c + 5c - 10 = 35
10c^2 - 15c - 10 - 35 = 0
10c^2 - 15c - 45 = 0
simplify, divide by 5
2c^2 - 3c - 9 = 0
Factors to
(2c+3)(c-3) = 0
the positive solution is all we want here
c = 3 yrs is Carl's present age
then
10(3) = 30 yrs is M's present age